Tax. Determine the SWL.
D: P = 440 - 2Q
S: P = 40 + 2Q
Tax: P = 80Q
1
Expert's answer
2016-05-20T12:40:03-0400
If demand and supply functions are:
D: P = 440 - 2Q
S: P = 40 + 2Q
And tax function is P = 80Q, then:
quantity demanded is 440 - 2Q = 80Q,
82Q = 440,
Q = 5.37 units.
quantity supplied is 40 + 2Q = 80Q,
78Q = 40,
Q = 0.51 units.
So, the SWL is 5.37 - 0.51 = 4.86 units.